OpenStudy (anonymous):

What polynomial has roots of –4, –1, and 6 ?

OpenStudy (nenadmatematika):

(x+4)(x+1)(x-6)

OpenStudy (saifoo.khan):

x = -4, x = -1 x = 6 (x + 4) = 0 (x + 1) = 0 (x - 6) = 0 (x+4)(x+1)(x-6) = 0 Now multiply the terms! :D

OpenStudy (anonymous):

x3 – x2 – 26x – 24 x3 – 9x2 – 22x + 24 x3 + x2 – 26x + 24 x3 + 9x2 + 14x – 24

OpenStudy (anonymous):

lamoo sorry you need these

OpenStudy (jamesj):

If a polynomial P(x) has r as a root, then we can write \[ P(x) = (x-r)Q(x) \] For some other polynomial Q(x). Hence generalizing, if \( r_1, r_2 \) and \( r_3 \) are roots, then \[ P(x) = (x-r_1)(x-r_2)(x-r_3)Q(x) \] In the solution just given to you, Q(x) = 1.

OpenStudy (precal):

multiply the solution out.